# Stanford CA 94305, USA

### v2

@@ 1 Introduction
We will assume that the method for getting a confidence interval involves randomizing the points at which $f$ is evaluated, via a pseudo-random number generator. This depends on using a well tested pseudo-random number generator. The big crush of [53] provides a thorough test. It is also possible to use a Bayesian model in which $f$ itself is random, but there is no counterpart to the big crush for that approach.
@@
A second tradeoff can arise. We may have certainty or just high confidence that $-\Delta_{n}\leqslant\hat{\mu}_{n}-\mu\leqslant\Delta_{n}$ where we also know that $|\hat{\mu}_{n}-\mu|/\Delta_{n}$ converges to zero, either deterministically or probabilistically. Then our resulting estimate $\Delta_{n}$ of the size of the error will for large enough $n$ be a signficant exaggeration, by an unknown factor. The bounds can be very conservative.
Since this article is to appear in an MCQMC proceedings, we assume some background knowledge that would take too much space to include here. For readers new to those topics, here are some bibliographic references that also include some history. Those topics include discrepancy [11], the ANOVA decomposition [69, Appendix A], weighted Hilbert spaces [22], digital QMC constructions [23] (such as Sobol’ sequences and Faure sequences) integration lattices [79] and their randomizations [49]. Some further references are given in context.
@@ 2 Notation and basic estimates of $\mu$
This paper surveys many different results from multiple literatures. As a result, it is necessary for some symbols to be used with more than one meaning. The specific use cases are different enough to be clear from context.
@@
for $d\geqslant 1$. Formulating a given problem this way may require extensive use of transformations such as those in [17] in order to handle domains other than $[0,1]^{d}$ and distributions other than the uniform one. Section 8 points to some new work using normalizing flows to augment those transformations. For very smooth $f$ and small $d$, classical quadrature methods from [15] are quite effective at estimating $\mu$.
@@
is the local discrepancy at ${\bm{a}}$. The QMC counterpart to the law of large numbers is that if $f$ is Riemann integrable and $D_{n}^{*}\to 0$ as $n\to\infty$, then $\hat{\mu}_{n}\to\mu$. See [63] which also mentions a converse that if $\hat{\mu}_{n}\to\mu$ whenever $D_{n}^{*}\to 0$, then $f$ must be Riemann integrable. We will see below that if $f$ is of bounded variation in the sense of Hardy and Krause (written $f\in\mathrm{BVHK}$) then $\varepsilon_{n}=O(D_{n}^{*})$. Since it is possible to find points with $D_{n}^{*}=O(n^{-1+\epsilon})$ for any $\epsilon>0$, we can get asymptotically lower errors from QMC than the RMSE of MC. The $\epsilon$ hides powers of $\log(n)$. Those are present in some worst cases, but they do not seem to appear in applications [70].
@@ 3 Uncertainty quantification
Equation (2) falls short of providing an uncertainty quantification because we ordinarily do not know either of the factors in it. It is expensive to compute $D_{n}^{*}$. The commonly used algorithm of [25] costs $O(n^{1+d/2})$. In the time it would take us to compute $D_{n}^{*}$ for QMC we could get $N=\Omega(n^{1+d/2})$ points in MC. Then MC would get an RMSE of $O(n^{-1/2-d/4})$ in the time that QMC using known $D_{n}^{*}$ would get an error of $O(n^{-1+\epsilon})$. We could in principle precompute $D_{n}^{*}$ for a number of point sets, but then we would face the second problem: $V_{\mathrm{HK}}(f)$ is extremely hard to compute, far worse than $\mu$ itself. For smooth enough $f$, it is a sum of $12^{d}-$ 1 integrals of the absolute values of some mixed partial derivatives of $f$ [67]. For less smooth $f$, the problem is even harder. As a result, (2) serves to show us that QMC can be much better than MC, but it does not generally let us verify that that has happened for given $f$ and $n$.
@@
The Chebyshev confidence interval is not commonly used. Even when we have a bound for $\sigma$, the interval can be very conservative. As mentioned above, to get 99% confidence, we can use $\lambda=10$. A Gaussian random variable only has probability about $1.5\times 10^{-23}$ to be $10$ or more standard deviations from its mean. The error probability bound of $0.01$ is quite loose (despite the fact that it came from an inequality that is tight).
@@
$\lim_{n\to\infty}\Pr\Bigl{(}\frac{\varepsilon_{n}}{\sigma/\sqrt{n}}\leqslant z\Bigr{)}=\Phi(z)$ (6)
@@
The usual statistical confidence interval is a bit wider than the one in (8) because it makes an adjustment for the sampling uncertainty in $s^{2}$. The interval in equation (8) is not even exact when $f({\bm{x}})$ has a Gaussian distribution. We may correct this using
@@
where $t^{1-\alpha/2}_{(n-1)}$ is the $21-\alpha/$ 2 quantile of Student’s $t$ distibution on $n-1$ degrees of freedom. The correction for degrees of freedom makes a meaningful difference when $n$ is small which is valuable for RQMC as we will see in Section 5 . The coverage is exactly $1-\alpha$ if $f({\bm{x}}_{i})$ are Gaussian (and $n\geqslant 2$) and the limit (9) holds if $f({\bm{x}})$ has finite variance.
@@
Then their Corollary 2 shows that a Borel set constructed based on $Y_{1},\dots,Y_{N}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}F$ that contains $\mu(F)$ with probability at least $1-\alpha$ also contains any other $m\in{\mathbb{R}}$ with probability at least $1-\alpha$. More precisely: we can get a confidence set, but not a useful one. They allow $N$ to be random so long as $\Pr(N<\infty)=1$.
@@
Suppose that independent $f({\bm{x}}_{i})$ satisfies known finite upper and lower bounds. Hoeffding’s inequality [46] then provides confidence intervals for $(1/n)\sum_{i=1}^{n}f({\bm{x}}_{i})$ in terms of those bounds. In the MC context it is natural to assume that the bounds are the same for all $i=1,\dots,n$. Because the bounds are known we can make a linear adjustment to $f$ and then without loss of generality we may then suppose that $0\leqslant f({\bm{x}})\leqslant 1$ for all ${\bm{x}}\in[0,1]^{d}$. Then for $t>0$, Hoeffding’s inequality gives
@@
There is ongoing work in forming confidence intervals that have guaranteed coverage for bounded random variables while not being too conservatively wide. See [4] and [87]. Some of this work allows for dependence among the values of $f({\bm{x}}_{i})$. Some of it provides confidence statements that are always valid in that they produce intervals $[a_{n},b_{n}]$ with $\Pr(a_{n}\leqslant\mu\leqslant b_{n},\forall n\geqslant 1)\geqslant 1-\alpha$. A sequence of intervals that is always valid has to be wider than if we only require validity at one $n$.
@@ 4 Certificates and bracketing
A certificate is a (finite) computable number that is mathematically guaranteed to be no smaller than $\varepsilon_{n}$. This concept is widely used in convex optimization [7]. When the goal is to minimize a convex function $f$ over ${\bm{x}}\in C$ for a convex set $C$, we can be sure that the minimum is no larger than $\min(f({\bm{x}}_{1}),\dots,f({\bm{x}}_{n}))$ which becomes known once $f$ has been evaluated at points ${\bm{x}}_{i}$ for $i=1,\dots,n$. Because $f$ does not go below any of its supporting hyperplanes the minimum must not be below the function value where any set of those hyperplanes intersect. That can provide a computable lower bound for the minimum of $f$. The user then has certainty that the answer lies between two computable numbers $a$ and $b$. If we take $(a+b)/2$ as the estimate then we know that the error is no larger than $(b-a)/2$.
@@
For smooth enough integrands, these error bounds from bracketing inequalities are far wider than necessary [15, Chapter 2]. If $f^{\prime}$ is integrable, then the left and right endpoint rules have an error that is asymptotically $O(1/n)$. If we average them then we get the trapezoid rule with error $O(1/n^{2})$ when $f^{\prime\prime}$ is continuous.
The same issue arises when we bracket the integrand between the midpoint and trapezoid rules. If $f^{\prime\prime}$ is continuous, then $\hat{\mu}_{\mathrm{Mid}}=\mu-f^{\prime\prime}(z_{\mathrm{M}})/(24n^{2})$ and $\hat{\mu}_{\mathrm{Trap}}=\mu+f^{\prime\prime}(z_{\mathrm{T}})/(12n^{2})$ for points $z_{\mathrm{M}},z_{\mathrm{T}}\in(0,1)$ These opposite signed errors can be largely canceled by Simpson’s rule. It satisfies $\hat{\mu}_{\mathrm{Simp}}=(2\hat{\mu}_{\mathrm{Mid}}+\hat{\mu}_{\mathrm{Trap}})/3$. If $f^{(4)}$ is continuous on $[0,1]$ then by equation (2.2.6) of [15]
@@
There is a discussion of practical error estimation in [15, Chapter 4.9]. One of the most applicable methods is to get two estimates, such as $\hat{\mu}_{n}$ and $\hat{\mu}_{2n}$. Then $|\hat{\mu}_{2n}-\hat{\mu}_{n}|$ is a rough estimate of the error in $\hat{\mu}_{n}$. It is also likely to be an overestimate of the error in $\hat{\mu}_{2n}$ which will usually be a better estimate than $\hat{\mu}_{n}$.
@@
The situation is very different for multivariate problems. Then there is a curse of dimension from [6]. Here is a sketch based on [24]. Suppose that the partial derivative of $f$ taken $a_{j}\geqslant 0$ times with respect to component $x_{j}$ is never larger than some $M\in(0,\infty)$ for any ${\bm{x}}\in[0,1]^{d}$ when $\sum_{j=1}^{d}a_{j}\leqslant r$. Then there still exists $B>0$ such that for any rule of the form $\sum_{i=1}^{n}w_{i}f({\bm{x}}_{i})$ for $w_{i}\in{\mathbb{R}}$ and ${\bm{x}}_{i}\in[0,1]^{d}$ there is $f$ with $|\hat{\mu}(f)-\mu(f)|>Bn^{-r/d}$. For randomized points ${\bm{x}}_{i}$, the RMSE cannot be below $B^{\prime}n^{-1/2-r/d}$ for some $B^{\prime}>0$. For large $d$, we cannot simply ignore the integration error. It is also not easy to use methods tuned for $r\geqslant 3$. Dimov [24] describes them as ‘sophisticated’.
@@
If we use ‘increasing’ as a less precise term but more vivid term for ‘nondecreasing’, then we may say that $f$ is increasing in every variable, the amount by which it is increasing in any variable is increasing in any other variable, that amount in turn is increasing in any third variable, and so on. This is a strict requirement. It is satisfied if $f$ is the cumulative distribution function (CDF) of some random random vector ${\bm{X}}$, i.e., $f({\bm{x}})=\Pr({\bm{X}}\leqslant{\bm{x}}\ \text{componentwise})$. As a partial converse, if $f$ is also right continuous, then
@@
holds for some probability measure $\nu$ and $\lambda\geqslant 0$ [1].
@@
To get a certificate we can use $n$ points that have NNLD and $n$ points that have NPLD. Gabai [32] showed that the Hammersley sequence in $[0,1]^{2}$ has NNLD. That is generalized in [33] to any digital net in $[0,1]^{d}$ where all of the generator matrices are permutation matrices. They also show that some rank one lattices have NNLD. The proofs of those NNLD results are based on the theory of associated random variables from reliability theory [30].
@@
The greater challenge than finding NNLD and NPLD points is the requirement that $f$ be completely monotone. It can be weakened by finding a control variate function $g$ with $\int_{[0,1]^{d}}g({\bm{x}}){\mathrm{d}}{\bm{x}}=0$ such that $f+g$ is completely monotone, and then averaging $(f+g)({\bm{x}}_{i})$. This has not been explored in the literature. For some $f$ it would work to take $g({\bm{x}})=c\prod_{j=1}^{d}(x_{j}-1/2)$ for large enough $c>0$. Such a function $g$ is not ‘QMC friendly’ because its ANOVA decomposition has all of its variance in the highest-order $d$-dimensional term. As a result, for large $d$ we expect $(1/n)\sum_{i=1}^{n}g({\bm{x}}_{i})$ to converge slowly to $\mu$. For instance, $cg$ has a very large Sobolev norm in the usual unanchored spaces. Using (3), we get $\partial^{u}g=c\prod_{j\not\in u}(x_{j}-1/2)$ and then
$\|g\|_{{\bm{\gamma}}}^{2}$ $=\sum_{u\subseteq 1{:}^{2}d}\frac{1}{\gamma_{u}}\int_{[0,1]^{|u|}}\biggl{(}\int_{[0,1]^{d-|u|}}c\prod_{j\not\in u}(x_{j}-1/2){\mathrm{d}}{\bm{x}}_{-u}\biggr{)}^{2}{\mathrm{d}}{\bm{x}}_{u}$
@@
For $\eta=2+\varepsilon$ with $\varepsilon>0$, we get strong tractability. The number of function evaluations to get a QMC error for $cg$ below $\epsilon$ can be brought below $c(d!)^{1+\varepsilon/2}$ times a constant that depends on $\epsilon$ but does not depend on $d$. The resulting rate in $d$ is quite unfavorable because $g$ has a very high norm in the weighted space.

### v3

@@ 1 Introduction
We will assume that the method for getting a confidence interval involves randomizing the points at which $f$ is evaluated, via a pseudo-random number generator. This depends on using a well tested pseudo-random number generator. The big crush of [63] provides such a thorough test. It is also possible to use a Bayesian model in which $f$ itself is random. The most common choice is a Gaussian process model for $f$. For an entry point to that large literature, one can start with [9]. The Bayesian approach can produce very accurate estimates but there is no counterpart to the big crush for that approach, and so it does not as yet have a well tested objective way to estimate error.
@@
A second tradeoff can arise. We may have certainty or just high confidence that $-\Delta_{n}\leqslant\hat{\mu}_{n}-\mu\leqslant\Delta_{n}$ where we also know that $|\hat{\mu}_{n}-\mu|/\Delta_{n}$ converges to zero, either deterministically or probabilistically. Then our resulting estimate $\Delta_{n}$ of the size of the error will for large enough $n$ be a significant exaggeration, by an unknown factor. The bounds can be very conservative.
Since this article is to appear in an MCQMC proceedings, we assume some background knowledge that would take too much space to include here. For readers new to those topics, here are some bibliographic references that also include some history. Those topics include discrepancy [13], the ANOVA decomposition [79, Appendix A], weighted Hilbert spaces [27], digital QMC constructions [28] (such as Sobol’ sequences and Faure sequences) integration lattices [92] and their randomizations [57]. Some further references are given in context. Further background is given on confidence interval methods as those are not widely studied in the QMC literature.
@@ 2 Notation and basic estimates of $\mu$
This paper surveys many different results from multiple literatures. As a result, it is necessary for some symbols to be used with more than one meaning, even as some symbols common in the literature have been changed to reduce the number of conflicts here. The specific use cases are different enough to be clear from context.
@@
for $d\geqslant 1$. Formulating a given problem this way may require extensive use of transformations such as those in [21] in order to handle domains other than $[0,1]^{d}$ and distributions other than the uniform one. Section 8 points to some new work using normalizing flows to augment those transformations. For very smooth $f$ and small $d$, classical integration methods from [19] are quite effective at estimating $\mu$.
@@
is the local discrepancy at ${\bm{a}}$. The QMC counterpart to the law of large numbers is that if $f$ is Riemann integrable and $D_{n}^{*}\to 0$ as $n\to\infty$, then $\hat{\mu}_{n}\to\mu$. See [73] which also mentions a converse that if $\hat{\mu}_{n}\to\mu$ whenever $D_{n}^{*}\to 0$, then $f$ must be Riemann integrable. We will see below that if $f$ is of bounded variation in the sense of Hardy and Krause (written $f\in\mathrm{BVHK}$) then $\varepsilon_{n}=O(D_{n}^{*})$. Since it is possible to find points with $D_{n}^{*}=O(n^{-1+\epsilon})$ for any $\epsilon>0$, we $\hat{\mu}_{n}=O(n^{-1+\epsilon})$, and then we can get asymptotically lower errors from QMC than the RMSE of MC.
The $\epsilon$ in that rate for $\varepsilon_{n}$ hides powers of $\log(n)$. Those are present in some worst cases, but they do not seem to appear in applications [81]. If $f$ has a continuous mixed partial derivative taken once with respect to each of the $d$ components of ${\bm{x}}$, then we can show that $f\in\mathrm{BVHK}$ using a result from [37]. That dominating mixed partial derivative falls short of the complete $r$-fold differentiability that appears in the $O(n^{-r/d})$ result mentioned earlier, and the price we pay is in logarithmic factors.
@@ 3 Uncertainty quantification
Equation (2) falls short of providing an uncertainty quantification because we ordinarily do not know either of the factors in it. It is expensive to compute $D_{n}^{*}$. The commonly used algorithm of [30] costs $O(n^{1+d/2})$. Let us suppose that it costs $\Omega(n^{1+d/2})$. Then in the time it would take us to compute $D_{n}^{*}$ for QMC we could get $N=\Omega(n^{1+d/2})$ points in MC. Then MC would get an RMSE of $O(n^{-1/2-d/4})$ in the time that QMC using known $D_{n}^{*}$ would get an error of $O(n^{-1+\epsilon})$. We could in principle precompute $D_{n}^{*}$ for a number of point sets, but then we would face the second problem: $V_{\mathrm{HK}}(f)$ is extremely hard to compute, far worse than $\mu$ itself. For smooth enough $f$, it is a sum of $12^{d}-$ 1 integrals of the absolute values of some mixed partial derivatives of $f$ [77]. For less smooth $f$, the problem is even harder. As a result, (2) serves to show us that QMC can be much better than MC, but it does not generally let us verify that that has happened for given $f$ and $n$.
@@
The Chebyshev confidence interval is not commonly used. Even when we have a bound for $\sigma$, the interval can be very conservative. As mentioned above, to get 99% confidence, we can use $\lambda=10$. A Gaussian random variable only has probability about $1.5\times 10^{-23}$ to be $10$ or more standard deviations from its mean. Here the error probability bound of $0.01$ is quite loose, despite the fact that it came from an inequality that is tight.
@@
$\lim_{n\to\infty}\Pr\Bigl{(}\frac{\hat{\mu}_{n}-\mu}{\sigma/\sqrt{n}}\leqslant z\Bigr{)}=\Phi(z)$ (6)
@@
The usual statistical confidence interval is a bit wider than the one in (8) because it makes an adjustment for the sampling uncertainty in $s^{2}$. The interval in equation (8) is not exact, even when $f({\bm{x}})$ has a Gaussian distribution. We may correct this using
@@
where $t^{1-\alpha/2}_{(n-1)}$ is the $21-\alpha/$ 2 quantile of Student’s $t$ distribution on $n-1$ degrees of freedom. The correction for degrees of freedom makes a meaningful difference when $n$ is small which is valuable for RQMC as we will see in Section 5 . The coverage is exactly $1-\alpha$ if $f({\bm{x}}_{i})$ are Gaussian (and $n\geqslant 2$) and the limit (9) holds if $f({\bm{x}})$ has finite variance.
@@
Under these conditions, their Corollary 2 shows that a Borel set constructed based on $Y_{1},\dots,Y_{N}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}F$ that contains $\mu(F)$ with probability at least $1-\alpha$ also contains any other $m\in{\mathbb{R}}$ with probability at least $1-\alpha$. More precisely: we can get a confidence set, but not a useful one. They allow $N$ to be random so long as $\Pr(N<\infty)=1$.
@@
Suppose that independent $f({\bm{x}}_{i})$ satisfies known finite upper and lower bounds. Hoeffding’s inequality [52] then provides confidence intervals for $\mu$ in terms of those bounds. In the MC context it is natural to assume that the bounds are the same for all $i=1,\dots,n$. Because the bounds are known we can make a linear adjustment to $f$ and then without loss of generality we may then suppose that $0\leqslant f({\bm{x}})\leqslant 1$ for all ${\bm{x}}\in[0,1]^{d}$. Then for $t>0$, Hoeffding’s inequality gives
@@
There is ongoing work in forming confidence intervals that have guaranteed coverage for bounded random variables and finite $n$ while being asymptotically as narrow as possible. See [4] and [101]. Some of this work allows for dependence among the values of $f({\bm{x}}_{i})$. Some of it provides confidence statements that are always valid in that they produce intervals $[a_{n},b_{n}]$ with $\Pr(a_{n}\leqslant\mu\leqslant b_{n},\forall n\geqslant 1)\geqslant 1-\alpha$. These confidence sequences have to be wider at each $n$ than if we only require validity at one value of $n$.
When $0\leqslant f({\bm{x}})\leqslant 1$ holds then this also holds for an average of $n$ evaluations of $f$ at some RQMC points. Jain et al. [53] study confidence interval widths using the empirical Bernstein and related betting methods from [101] for RQMC using $R$ random replicates of $n$ RQMC points. They impose a budget constraint that $nR=N$. If $f$ is regular enough for RQMC to be very effective, then the optimal $n$ grows only very slowly as $N\to\infty$. The resulting confidence intervals narrow at a faster rate in $N$ than the ones based on Monte Carlo do, but they narrow at a slower rate in $N$ than the RQMC standard deviations do.
@@ 4 Certificates and bracketing
A certificate is a (finite) computable number that is mathematically guaranteed to be no smaller than $\varepsilon_{n}$. This concept is widely used in convex optimization [8]. There the goal is to minimize a convex function $f$ over ${\bm{x}}\in C$ for a convex set $C$. Once $f$ has been evaluated at points ${\bm{x}}_{i}$ for $i=1,\dots,n$, we can be sure that the minimum is no larger than $\min(f({\bm{x}}_{1}),\dots,f({\bm{x}}_{n}))$. This upper bound does not require convexity of $f$. If $f$ is convex, then with enough data, we can get a lower bound. A convex function cannot go below any of its supporting hyperplanes. Then $f$ has to be above the pointwise maximum of all its supporting hyperplanes that the algorithm has encountered. Now the minimum value of that pointwise maximum provides a computable lower bound for the minimum of $f$. See Chapter 5 of [8] on duality. The user then has certainty that the answer lies between two computable numbers $a$ and $b$. If we take $(a+b)/2$ as the estimate then we know that the error is no larger than $(b-a)/2$.
@@
For smooth enough integrands, these error bounds from bracketing inequalities are far wider than necessary [19, Chapter 2]. If $f^{\prime}$ is integrable, then the left and right endpoint rules have an error that is asymptotically $O(1/n)$. If we average them then we get the trapezoid rule with error $O(1/n^{2})$ if $f^{\prime}\in L_{2}[0,1]$ [18].
The same issue arises when we bracket the integrand between the midpoint and trapezoid rules. If $f^{\prime\prime}$ is continuous, then $\hat{\mu}_{\mathrm{Mid}}=\mu-f^{\prime\prime}(z_{\mathrm{M}})/(24n^{2})$ and $\hat{\mu}_{\mathrm{Trap}}=\mu+f^{\prime\prime}(z_{\mathrm{T}})/(12n^{2})$ for points $z_{\mathrm{M}},z_{\mathrm{T}}\in(0,1)$. These oppositely signed errors can be largely canceled by Simpson’s rule. It satisfies $\hat{\mu}_{\mathrm{Simp}}=(2\hat{\mu}_{\mathrm{Mid}}+\hat{\mu}_{\mathrm{Trap}})/3$. If $f^{(4)}$ is continuous on $[0,1]$ then by equation (2.2.6) of [19]
@@
There is a discussion of practical error estimation in [19, Chapter 4.9]. Many of them require special conditions on $f$. One of the most applicable methods is to get two estimates, such as $\hat{\mu}_{n}$ and $\hat{\mu}_{2n}$. Then $|\hat{\mu}_{2n}-\hat{\mu}_{n}|$ is a rough estimate of the error in $\hat{\mu}_{n}$. It is also likely to be an overestimate of the error in $\hat{\mu}_{2n}$ which will usually be a better estimate than $\hat{\mu}_{n}$. Despite its wide use, this method can clearly give an unreliable estimate. There is an analysis in [68] showing how using the difference between two different integration rules can be an unreliable way to quantify uncertainty.
@@
The situation is very different for multivariate problems. Then there is a curse of dimension from [6]. Here is a sketch based on [29]. Suppose that the absolute value of the partial derivative of $f$ taken $a_{j}\geqslant 0$ times with respect to component $x_{j}$ is never larger than some $M\in(0,\infty)$ for any ${\bm{x}}\in[0,1]^{d}$ when $\sum_{j=1}^{d}a_{j}\leqslant r$. Then there still exists $B>0$ such that for any rule of the form $\sum_{i=1}^{n}w_{i}f({\bm{x}}_{i})$ for $w_{i}\in{\mathbb{R}}$ and ${\bm{x}}_{i}\in[0,1]^{d}$ there is $f$ with $|\hat{\mu}(f)-\mu(f)|>Bn^{-r/d}$. For randomized points ${\bm{x}}_{i}$, the RMSE cannot be below $B^{\prime}n^{-1/2-r/d}$ for some $B^{\prime}>0$. For large $d$, we cannot simply ignore the integration error. It is also not easy to use methods tuned for $r\geqslant 3$. Dimov [29] describes them as ‘sophisticated’.
@@
If we use ‘increasing’ as a less precise but more vivid term for ‘nondecreasing’, then we may say that $f$ is increasing in every variable, the amount by which it is increasing in any variable is increasing in any other variable, that amount in turn is increasing in any third variable, and so on. This is a strict requirement. It is satisfied if $f$ is the cumulative distribution function (CDF) of some random random vector ${\bm{X}}$, i.e., $f({\bm{x}})=\Pr({\bm{X}}\leqslant{\bm{x}}\ \text{componentwise})$. As a partial converse, if $f$ is also right continuous, then
@@
holds for some probability measure $\nu$ and some $\lambda\geqslant 0$ [1].
@@
To get a certificate we can use $n$ points that have NNLD and $n$ points that have NPLD. Gabai [38] showed that the Hammersley sequence in $[0,1]^{2}$ (in any base $b\geq 2$) has NNLD. That is generalized in [39] to any digital net in $[0,1]^{d}$ where all of the generator matrices are permutation matrices. They also show that some rank one lattices have NNLD. The proofs of those NNLD results are based on the theory of associated random variables from reliability theory [35].
@@
The greater challenge than finding NNLD and NPLD points is the requirement that $f$ be completely monotone. It can be weakened by finding a control variate function $g$ with $\int_{[0,1]^{d}}g({\bm{x}}){\mathrm{d}}{\bm{x}}=0$ such that $f+g$ is completely monotone, and then averaging $(f+g)({\bm{x}}_{i})$. This has not been explored in the literature. For some $f$ it would work to take $g({\bm{x}})=c\prod_{j=1}^{d}(x_{j}-1/2)$ for large enough $c>0$. Such a function $g$ is not ‘QMC friendly’ because its ANOVA decomposition has all of its variance in the highest-order $d$-dimensional term. As a result, for large $d$ we expect $(1/n)\sum_{i=1}^{n}g({\bm{x}}_{i})$ to converge slowly to $\mu$. For instance, $g$ has a very large Sobolev norm in the usual unanchored spaces. Using (3) for $\gamma_{j}=j^{-\eta}$ we get $\partial^{u}g=c\prod_{j\not\in u}(x_{j}-1/2)$ and then
$\|g\|_{{\bm{\gamma}}}^{2}$ $=c^{2}\sum_{u\subseteq 1{:}d}\frac{1}{\gamma_{u}}\int_{[0,1]^{|u|}}\biggl{(}\int_{[0,1]^{d-|u|}}\prod_{j\not\in u}(x_{j}-1/2){\mathrm{d}}{\bm{x}}_{-u}\biggr{)}^{2}{\mathrm{d}}{\bm{x}}_{u}$
@@
For $\eta=2+\varepsilon$ with $\varepsilon>0$, we get strong tractability. The number of function evaluations to get a QMC error for $g$ below $\epsilon$ can be brought below $c(d!)^{1+\varepsilon/2}$ times a constant that depends on $\epsilon$ but does not depend on $d$. The resulting rate in $d$ is quite unfavorable because $g$ has a very high norm in the weighted space.